Latisha puts a fence around her back yard. She finds that she uses 15 more feet for each long side of the yard than she does for each width. If she used 70ft of fencing, how long is each side of her yard(answer with a length and a width).
Found 2 solutions by stanbon, monika_p:
Answer by stanbon(75887)


Latisha puts a fence around her back yard. She finds that she uses 15 more feet for each long side of the yard than she does for each width. If she used 70ft of fencing, how long is each side of her yard(answer with a length and a width).
----------------
Let the width be "x",
then the length will be "x+15".
----------------
Perimeter: P = 2(L+W)
70 = 2(x + x+15)
2x+15 = 35
2x = 20
x = 10 ft. (width)
x+15 = 15 ft (length)
======================
Cheers,
Stan H.






Answer by monika_p(71)

lets x be widht of the yard, then lenght of the yard y=x+15 and perimeter P=70
formula for perimeter of rectangle is P=2x+2y,then
70=2*(x+15)+2x
70=4x+30
x=10
y=25